OpenStudy (anonymous):

i have 136 people going on a zoo feild trip and 13$ for adult and 8$ for child cost them 1,163 . How many adults went and how many children went? I got -8x -8 = 1088 and 13x+8y=1163 5x=75 adults but cant figure out how they got the 1088 on that problem

OpenStudy (anonymous):

you have a good start; two variables means you need two equations. Set x to adults, and y to children. \[x+y=136\] \[13x+8y=1163\] Substitute the first equation into the second: \[x = 136-y\] \[13(136-y)+8y=1163\] \[1768-13y+8y=1163\] \[5y = 605\] \[y = 121\] \[x = 136 - 121 = 15\]

OpenStudy (anonymous):

Damn sorry for double post! You should definitely understand the process now! haha

OpenStudy (anonymous):

thanks alot for help

OpenStudy (anonymous):

No problem